OpenStudy (anonymous):

2(5)+3y=9 When you answer can you write out how u do it?:)

OpenStudy (phoenixfire):

Are you solving for Y?

OpenStudy (anonymous):

Yes

OpenStudy (phoenixfire):

Do the multiplication. 2(5)=?

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

10

OpenStudy (phoenixfire):

Right, so you have \[10+3y=9\]You want to isolate terms with 'y' in them. So subtract 10 from both sides. What do you get?

OpenStudy (anonymous):

Idk

OpenStudy (phoenixfire):

\[3y+10-10=9-10\] Can you simplify that?